# algorithmic

### v1

@@ 4.1.2 Online classification with periodic drift
We can improve the performance of RL[K]-PR if the number of hypotheses $K$ increases, and if we vary the changepoint probability threshold $\pi$, as shown in Figure 6. However, even then the performance of this method still does not match our method.
Figure 6: Accuracy of predictions for RL[1]-PR as a function of the number of hypothesis and the prior probability of a changepoint $\pi$. The black dotted line is the performance of RL[1]-OUPR* reported in Figure 5.
@@ 4.1.3 Online classification with drift and jumps
Figure 8: Accuracy of predictions for RL[K]-PR as a function of the number of hypotheses (K) and the probability of a changepoint $\pi$. The black dotted line is the performance of RL[1]-OUPR* reported in Figure 7.
To explain this behaviour, Figure 8 shows the performance of RL[K]-PR as a function of number of hypotheses and prior probability of a changepoint $\pi$. We observe that up to three hypotheses, the lowest misclassification error of RL[K]-PR is higher than that of RL[1]-OUPR*, which only considers one hypothesis. However, as we increase the number of hypotheses, the best performance for RL[K]-PR obtains a lower misclassification rate than RL[1]-OUPR*. This is in contrast to the results in Figure 5. Here, we see that with more hypotheses RL[K]-PR outperforms our new method at the expense of being more memory intensive.
@@ 4.3.1 Autoregression with dependence across the segments
We compare the performance of the three methods in the previous subsection. For this experiment, we employ a probability of a changepoint $\pi=0.01$. Since this dataset has dependence of the parameters across segments, we allow for the choice of (M.1: likelihood) to be influenced by the choice of (M.2: auxvar), i.e., the conditional expectation (5) can be written as $h({\bm{\theta}}_{t};\psi_{t},{\bm{x}}_{t})$. For this experiment, we take (M.2: auxvar) to be RL and our choice of (M.1: likelihood) becomes
@@
with ${\bm{h}}({\bm{x}}_{1:t},r_{t})=[1,\Delta,\Delta^{2}]$, $\Delta=(x_{t}-x_{r_{t}})$, and $x_{r_{t}}\geq x_{t}$. Intuitively this represents a quadratic curve fit to the beginning $x_{r_{t}}$ and end points $x_{t}$ of the current segment. Given the form of (M.1: likelihood) in (40), here we do not consider C-ACI nor CPP-OU. Instead, we use runlength with moment-matching prior reset, i.e., RL-MMPR (see Table 3) which was designed for segmentation with dependence.
@@ 4.3.2 Non-stationary heavy-tailed regression with DA[inf]
It is well-known that the combination RL-PR is sensitive to outliers if the choice of (M.1: likelihood) is misspecified, since an observation that is “unusual” may trigger a changepoint unnecessarily. As a consequence, various works have proposed outlier-robust variants to the RL[inf]-PR for segmentation (Knoblauch et al., 2018; Fearnhead & Rigaill, 2019; Altamirano et al., 2023; Sellier & Dellaportas, 2023) and for filtering (Reimann, 2024). In what follows, we show how we can easily accomodate robust methods into the BONE framework by changing the way we compute the likelihood and/or posterior. In particular, we consider the WoLF-IMQ method of Duran-Martin et al. (2024).333 We set the soft threshold value to 4, representing four standard deviations of tolerance before declaring an outlier. We use WoLF-IMQ because it is a provably robust algorithm and it is a straightforward modification of the linear Gaussian posterior update equations. We denote RL[inf]-PR with (A.1: posterior) taken to be LG as LG+RL[inf]-PR and RL[inf]-PR with (A.1: posterior) taken to be WoLF-IMQ as WoLF+RL[inf]-PR*.
@@
To process this data, our choice of (M.1: likelihood) is $h({\bm{\theta}}_{t},{\bm{x}}_{t})={\bm{\theta}}_{t}^{\intercal}\,{\bm{x}}_{t}$.
@@ 5 Conclusions
We introduced a unified Bayesian framework to perform online predictions in non-stationary environments, and showed how it covers many prior works. We also used our framework to design a new method, RL[1]-OUPR*, which is suited to tackle prediction problems when the observations exhibit both abrupt and gradual changes. We hope to explore other novel variants and applications in future work.
6 Aknowledgments
We thank Matias Altamirano, François-Xavier Briol, Patrick Chang, Alex Galashov, Matt Jones, Jeremias Knoblauch, and Radford Neal for insightful comments and discussions.
@@ Appendix A Worked examples for BONE methods
In this section, we provide a detailed calculation of $\nu_{t}(\psi_{t})$ for some choices of $\psi_{t}$. We consider a choice of (M.1: likelihood) to be linear Gaussian with known observation variance $\mathbf{R}_{t}$, i.e.,
@@ A.1.1 Unbounded number of hypotheses RL[inf]-PR
For a choice of (M.1: likelihood) given by (42) and a choice of (M.3: prior) given by (16), the posterior predictive (46) takes the form.
@@ A.3 Runlength with OU dynamics and prior reset (RL[1]-OUPR*)
In essence, RL[1]-OUPR* follows the logic behind RL[1]-PR introduced in Section A.1 with $K=1$ hypothesis and different choice of (M.3: prior). To derive the algorithm for RL[1]-OUPR* at time $t>1$, suppose $r_{t-1}$ is available (the only hypothesis we track). Denote by $r_{t}^{(1)}$ the hypothesis of a runlength increase, i.e., $r_{t}=r_{t-1}+1$ and denote by $r_{t}^{(0)}$ the hypothesis of a runlenght reset, i.e., $r_{t}=0$. The probability of a runlength increase under a single hypothesis takes the form
$\nu_{t}(r_{t})$ $=p(r_{t}^{(1)}\,|\,\mathcal{D}_{1:t})$ (54)
@@
$=\frac{p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t})\,p(r_{t-1},\mathcal{D}_{1:t-1})\,(1-\pi)}{p({\bm{y}}_{t}\,|\,r_{t}^{(0)},{\bm{x}}_{t},\mathcal{D}_{1:t})\,p(r_{t-1},\mathcal{D}_{1:t-1})\,\pi+p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t})\,p(r_{t-1},\mathcal{D}_{1:t-1})\,(1-\pi)}$
$=\frac{p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t})\,(1-\pi)}{p({\bm{y}}_{t}\,|\,r_{t}^{(0)},{\bm{x}}_{t},\mathcal{D}_{1:t})\,\pi+p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t})\,(1-\pi)}.$
Next, we use $\nu_{t}(r_{t})$ to decide whether to update our parameters or reset them according to a prior belief according to some threshold $\epsilon$. This implements our choice of (M.3: prior) given in (17) and (18). Because we maintain a single hypothesis, the weight at the end of the update step is set to $1$. Algorithm 4 shows an update step for RL[1]-OUPR* under the choice of (M.1: likelihood) given by (42).
@@ Appendix B Algorithms
0:  $p({\bm{y}}\,|\,{\bm{\theta}},{\bm{x}})={\cal N}({\bm{y}}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}},\mathbf{R}_{t})$ // Choice of (M.1: likelihood)
@@
8:     $\lambda_{t}({\bm{\theta}}_{t};\,r_{t}^{(k)})\propto\tau_{t}({\bm{\theta}}_{t};r_{t}^{(k)})\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}}_{t},\mathbf{R}_{t})\propto{\cal N}({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu}_{t}^{(k)},\boldsymbol{\Sigma}_{t}^{(k)})$ // following (24)
@@
0:  $p({\bm{y}}\,|\,{\bm{\theta}},{\bm{x}})={\cal N}({\bm{y}}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}},\mathbf{R}_{t})$ // Choice of (M.1: likelihood)
@@
8:     $\lambda_{t}({\bm{\theta}}_{t};\,r_{t}^{(k)})\propto\tau_{t}({\bm{\theta}}_{t};r_{t}^{(k)})\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}}_{t},\mathbf{R}_{t})\propto{\cal N}({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu}_{t}^{(k)},\boldsymbol{\Sigma}_{t}^{(k)})$ // following (24)
@@
0:  $p({\bm{y}}\,|\,{\bm{\theta}},{\bm{x}})={\cal N}({\bm{y}}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}},\mathbf{R}_{t})$ // Choice of (M.1: likelihood)
@@
16:  $\lambda_{t}({\bm{\theta}}_{t};\,r_{t})\propto{\cal N}({\bm{\theta}}_{t}\,|\,\bar{\boldsymbol{\mu}}_{t},\bar{\boldsymbol{\Sigma}}_{t})\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}}_{t},\mathbf{R}_{t})\propto{\cal N}({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t})$ // choice of (A.1: posterior)— via (24)

### v3

@@ 4.1.2 Online classification with periodic drift
We can improve the performance of RL[K]-PR if the number of hypotheses $K$ increases, and if we vary the changepoint probability threshold $\kappa$, as shown in Figure 8. However, even then the performance of this method still does not match our method.
Figure 8: Accuracy of predictions for RL[1]-PR as a function of the number of hypothesis and the prior probability of a changepoint $\kappa$. The black dotted line is the performance of RL[1]-OUPR* reported in Figure 7.
@@ 4.1.3 Online classification with drift and jumps
Figure 10: Accuracy of predictions for RL[K]-PR as a function of the number of hypotheses (K) and the probability of a changepoint $\kappa$. The black dotted line is the performance of RL[1]-OUPR* reported in Figure 9.
To explain this behaviour, Figure 10 shows the performance of RL[K]-PR as a function of number of hypotheses and prior probability of a changepoint $\kappa$. We observe that up to three hypotheses, the lowest misclassification error of RL[K]-PR is higher than that of RL[1]-OUPR*, which only considers one hypothesis. However, as we increase the number of hypotheses, the best performance for RL[K]-PR obtains a lower misclassification rate than RL[1]-OUPR*. This is in contrast to the results in Figure 5. Here, we see that with more hypotheses RL[K]-PR outperforms our new method at the expense of being more memory intensive.
@@ 4.3.1 Autoregression with dependence across the segments
We compare the performance of the three methods in the previous subsection. For this experiment, we employ a probability of a changepoint $\kappa=0.01$. Since this dataset has dependence of the parameters across segments, we allow for the choice of (M.1: model) to be influenced by the choice of (M.2: auxvar), i.e., our choice of model is given by $h({\bm{\theta}}_{t};\psi_{t},{\bm{x}}_{t})$. For this experiment, we take (M.2: auxvar) to be RL and our choice of (M.1: model) becomes
@@
with ${\bm{h}}({\bm{x}}_{1:t},r_{t})=[1,\Delta,\Delta^{2}]$, $\Delta=(x_{t}-x_{r_{t}})$, and $x_{r_{t}}\geq x_{t}$. Intuitively this represents a quadratic curve fit to the beginning $x_{r_{t}}$ and end points $x_{t}$ of the current segment. Given the form of (M.1: model) in (44), here we do not consider C-ACI nor CPP-OU. Instead, we use runlength with moment-matching prior reset, i.e., RL-MMPR (see Table 4) which was designed for segmentation with dependence.
@@ 4.3.2 Non-stationary heavy-tailed regression with DA[inf]
It is well-known that the combination RL-PR is sensitive to outliers if the choice of (M.1: model) is misspecified, since an observation that is “unusual” may trigger a changepoint unnecessarily. As a consequence, various works have proposed outlier-robust variants to the RL[inf]-PR for segmentation (Knoblauch et al., 2018; Fearnhead & Rigaill, 2019; Altamirano et al., 2023; Sellier & Dellaportas, 2023) and for filtering (Reimann, 2024). In what follows, we show how we can easily accomodate robust methods into the BONE framework by changing the way we compute the likelihood and/or posterior. In particular, we consider the WoLF-IMQ method of Duran-Martin et al. (2024). We use WoLF-IMQ because it is a provably robust algorithm and it is a straightforward modification of the linear Gaussian posterior update equations. We denote RL[inf]-PR with (A.1: posterior) taken to be LG as LG+RL[inf]-PR and RL[inf]-PR with (A.1: posterior) taken to be WoLF-IMQ as WoLF+RL[inf]-PR*.
@@
To process this data, our choice of (M.1: model) is $h({\bm{\theta}}_{t},{\bm{x}}_{t})={\bm{\theta}}_{t}^{\intercal}\,\phi({\bm{x}}_{t})$ with
$\ell({\bm{y}}_{t};{\bm{\theta}}_{t},{\bm{x}}_{t})=-W^{2}\big{(}{\bm{y}}_{t},h({\bm{\theta}}_{t},{\bm{x}}_{t})\big{)}\,\log{\cal N}({\bm{y}}_{t}\,|\,h({\bm{\theta}}_{t},{\bm{x}}_{t}),1.0),$ (46)
a weighted Gaussian log-likelihood and $W(u,z)=(1+\frac{(u-z)^{2}}{c^{2}})^{-1/2}$ the inverse multi-quadratic (IMQ) function with soft threshold value $c=4$, representing four standard deviations of tolerance to outliers. Here $u,z\in$.
@@ 5 Conclusions
We introduced a unified Bayesian framework to perform online predictions in non-stationary environments, and showed how it covers many prior works. We also used our framework to design a new method, RL[1]-OUPR*, which is suited to tackle prediction problems when the observations exhibit both abrupt and gradual changes. In future work, we aim to investigate other novel variants and applications to newer model architectures such as transformers and graph neural networks (Moreno-Pino et al., 2024; Arroyo et al., 2025).
Aknowledgments
We thank Matias Altamirano, François-Xavier Briol, Patrick Chang, Alex Galashov, Matt Jones, Jeremias Knoblauch, and Radford Neal for insightful comments and discussions on earlier versions of this paper.
@@ Appendix A Worked examples for BONE methods
In this section, we provide a detailed calculation of $\nu_{t}(\psi_{t})$ for some choices of $\psi_{t}$. We consider a choice of (M.1: model) to be linear Gaussian with known observation variance $\mathbf{R}_{t}$, i.e.,
@@ A.1.1 Unbounded number of hypotheses RL[inf]-PR
For a choice of (M.1: model) given by (47) and a choice of (M.3: prior) given by (18), the posterior predictive (51) takes the form.
@@ A.3 Runlength with OU dynamics and prior reset (RL[1]-OUPR*)
In essence, RL[1]-OUPR* follows the logic behind RL[1]-PR introduced in Section A.1 with $K=1$ hypothesis and different choice of (M.3: prior). To derive the algorithm for RL[1]-OUPR* at time $t>1$, suppose $r_{t-1}$ is available (the only hypothesis we track). Denote by $r_{t}^{(1)}$ the hypothesis of a runlength increase, i.e., $r_{t}=r_{t-1}+1$ and denote by $r_{t}^{(0)}$ the hypothesis of a runlenght reset, i.e., $r_{t}=0$. The probability of a runlength increase under a single hypothesis takes the form:
$\nu_{t}(r_{t}^{(1)})$ $=p(r_{t}^{(1)}\,|\,\mathcal{D}_{1:t})$ (59)
@@
$=\frac{p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,p(r_{t-1},\mathcal{D}_{1:t-1})\,(1-\kappa)}{p({\bm{y}}_{t}\,|\,r_{t}^{(0)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,p(r_{t-1},\mathcal{D}_{1:t-1})\,\kappa+p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,p(r_{t-1},\mathcal{D}_{1:t-1})\,(1-\kappa)}$
$=\frac{p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,(1-\kappa)}{p({\bm{y}}_{t}\,|\,r_{t}^{(0)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,\kappa+p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,(1-\kappa)}.$
where $\kappa=p(r_{t}\,|\,r_{t-1})$ with $r_{t}=0$ is the prior probability of a changepoint and and $1-\kappa=p(r_{t}\,|\,r_{t-1}$ with $r_{t}=r_{t-1}+1$ is the probability of continuation of the current segment.
Next, we use $\nu_{t}(r_{t})$ to decide whether to update our parameters or reset them according to a prior belief according to some threshold $\epsilon$. This implements our choice of (M.3: prior) given in (19) and (20). Because we maintain a single hypothesis, the weight at the end of the update step is set to $1$. Algorithm 4 shows an update step for RL[1]-OUPR* under the choice of (M.1: model) given by (47).
@@ Appendix B Algorithms
0:  $p({\bm{y}}\,|\,{\bm{\theta}},{\bm{x}})={\cal N}({\bm{y}}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}},\mathbf{R}_{t})$ // Choice of (M.1: model)
@@
8:     $q_{t}({\bm{\theta}}_{t};\,r_{t}^{(k)})\propto\tau_{t}({\bm{\theta}}_{t};r_{t}^{(k)})\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}}_{t},\mathbf{R}_{t})\propto{\cal N}({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu}_{t}^{(k)},\boldsymbol{\Sigma}_{t}^{(k)})$ // following (26)
@@
0:  $p({\bm{y}}\,|\,{\bm{\theta}},{\bm{x}})={\cal N}({\bm{y}}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}},\mathbf{R}_{t})$ // Choice of (M.1: model)
@@
8:     $q_{t}({\bm{\theta}}_{t};\,r_{t}^{(k)})\propto\tau_{t}({\bm{\theta}}_{t};r_{t}^{(k)})\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}}_{t},\mathbf{R}_{t})\propto{\cal N}({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu}_{t}^{(k)},\boldsymbol{\Sigma}_{t}^{(k)})$ // following (26)
@@
0:  $p({\bm{y}}\,|\,{\bm{\theta}},{\bm{x}})={\cal N}({\bm{y}}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}},\mathbf{R}_{t})$ // Choice of (M.1: model)
@@
16:  $q_{t}({\bm{\theta}}_{t};\,r_{t})\propto{\cal N}({\bm{\theta}}_{t}\,|\,\bar{\boldsymbol{\mu}}_{t},\bar{\boldsymbol{\Sigma}}_{t})\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}^{\intercal}{\bm{x}}_{t},\mathbf{R}_{t})\propto{\cal N}({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t})$ // choice of (A.1: posterior)— via (26)
@@
Appendix C Bayesian inference when the dynamics are known
If the dynamics of model parameters $p({\bm{\theta}}_{t+1},\,|\,{\bm{\theta}}_{t},\psi_{t+1})$ and the dynamics for auxiliary variable $p(\psi_{t+1}\,|\,\psi_{t})$ are known, then a full-Bayesian treatment of the prediction at time $t+1$ is given by predicting the expected model parameters and then performing inference.
Specifically, given the prior belief state $p(\psi_{t},{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t})$ and the new input ${\bm{x}}_{t+1}$, the predictive distribution over the output ${\bm{y}}_{t+1}$ is given by
$p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t})$ $=\sum_{\psi_{t+1}}\int_{{\bm{\theta}}_{t+1}}p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},{\bm{\theta}}_{t+1})p(\psi_{t+1},{\bm{\theta}}_{t+1}\,|\,\mathcal{D}_{1:t})\,\mathrm{d}{\bm{\theta}}_{t+1},$ (64)
where the parameter predictive distribution is
$p(\psi_{t+1},{\bm{\theta}}_{t+1}\,|\,\mathcal{D}_{1:t})$ $=\sum_{\psi_{t}}\int_{{\bm{\theta}}_{t}}p(\psi_{t+1},{\bm{\theta}}_{t+1}\,|\,\psi_{t},{\bm{\theta}}_{t})p(\psi_{t},{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t})\,\mathrm{d}{\bm{\theta}}_{t}.$ (65)
Here, the parameter dynamics factorises as
$p(\psi_{t+1},{\bm{\theta}}_{t+1}\,|\,\psi_{t},{\bm{\theta}}_{t})$ $=p(\psi_{t+1}\,|\,\psi_{t})p({\bm{\theta}}_{t+1}\,|\,{\bm{\theta}}_{t},\psi_{t+1}),$ (66)
and the prior belief state factorises as
$p(\psi_{t},{\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t})$ $=p(\psi_{t}\,|\,\mathcal{D}_{1:t})p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t}).$ (67)
Combining all the above equations, we can derive an expression for the posterior predicted mean
$\hat{{\bm{y}}}_{t+1}$ $=\mathbb{E}[{\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t}]$ (68)
$=\sum_{\psi_{t+1}}\sum_{\psi_{t}}p(\psi_{t+1}\,|\,\psi_{t})p(\psi_{t}\,|\,\mathcal{D}_{1:t})$
×$\qquad\times\int_{{\bm{\theta}}_{t+1}}\int_{{\bm{\theta}}_{t}}h({\bm{\theta}}_{t+1},{\bm{x}}_{t+1})p({\bm{\theta}}_{t+1}\,|\,{\bm{\theta}}_{t},\psi_{t+1})p({\bm{\theta}}_{t}\,|\,\psi_{t},\mathcal{D}_{1:t})\,\mathrm{d}{\bm{\theta}}_{t}\,\,\mathrm{d}{\bm{\theta}}_{t+1}.$
After making the prediction, we observe the outcome ${\bm{y}}_{t+1}$, and then update the belief state for the parameter ${\bm{\theta}}_{t+1}$ (for each value of $\boldsymbol{\psi}_{t+1}$) using
$p({\bm{\theta}}_{t+1}\,|\,\psi_{t+1},\mathcal{D}_{1:t+1})$ $=\frac{1}{Z_{t+1}^{\psi_{t+1}}}p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},{\bm{\theta}}_{t+1})p({\bm{\theta}}_{t+1}\,|\,\psi_{t+1},\mathcal{D}_{1:t}),$ (69)
$Z_{t+1}^{\psi_{t+1}}$ $=p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t},\psi_{t+1})=\int_{{\bm{\theta}}_{t+1}}p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},{\bm{\theta}}_{t+1})p({\bm{\theta}}_{t+1}\,|\,\psi_{t+1},\mathcal{D}_{1:t})\,\mathrm{d}{\bm{\theta}}_{t+1},$
$p({\bm{\theta}}_{t+1}\,|\,\psi_{t+1},\mathcal{D}_{1:t})$ $=\int_{{\bm{\theta}}_{t}}p({\bm{\theta}}_{t+1}\,|\,{\bm{\theta}}_{t},\psi_{t+1})p({\bm{\theta}}_{t}\,|\,\mathcal{D}_{1:t})\,\mathrm{d}{\bm{\theta}}_{t}.$
Similarly we update the belief state for the auxiliary variable using
$p(\psi_{t+1}\,|\,\mathcal{D}_{1:t+1})$ $=\frac{1}{Z_{t+1}}p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\psi_{t+1})p(\psi_{t+1}\,|\,\mathcal{D}_{1:t}),$ (70)
$Z_{t+1}$ $=p({\bm{y}}_{t+1}\,|\,{\bm{x}}_{t+1},\mathcal{D}_{1:t})=\sum_{\psi_{t+1}}Z_{t+1}^{\psi_{t+1}}p(\psi_{t+1}\,|\,\mathcal{D}_{1:t}),$
$p(\psi_{t+1}\,|\,\mathcal{D}_{1:t})$ $=\sum_{\psi_{t}}p(\psi_{t+1}\,|\,\psi_{t})\,p(\psi_{t}\,|\,\mathcal{D}_{1:t}).$
This full Bayesian treatment is different from BONE, because we do not assume well-specified dynamics. Thus, BONE makes predictions at time $t+1$ with only the posterior at time $t$, i.e., it does not require a predict step.